The game of dots is played by rolling a fair die and recieving $1 for each dot showing on the top face of the die. What cost should be set for each roll if the game is to be considered a fair game?
Thanks in advance for your help.

Answer by stanbon(75887)

The game of dots is played by rolling a fair die and recieving $1 for each dot showing on the top face of the die. What cost should be set for each roll if the game is to be considered a fair game?
-----------------------
The probability of each number value is 1/6.
The expected earnings on the game is:
(1/6)(1 + 2 + 3 + 4 + 5 + 6) = (1/6)*21 = 7/2 = 3.5
---------------
To make the game fair and charge should be $3.50
since the expected earnings is $3.50.
=============================================================
Cheers,
Stan H.
=============================================================